OpenStudy (anonymous):

Consider the diff eq dy/dt = ((e^y)*sin(t)^2)/y*sec(t))

OpenStudy (cruffo):

considering it...

OpenStudy (anonymous):

My work thus far is y*dy/dt = e^y sin^2(t)/sec(t) how do I switch e^y to the other side?

OpenStudy (anonymous):

do I multiply by ln(e^y)?

OpenStudy (cruffo):

is that \[\sin^2(t)\] or \[sin(t^2)\]

OpenStudy (anonymous):

sin^2(t)

OpenStudy (anonymous):

ydy/ e^y = sin^2 t cost dt

OpenStudy (cruffo):

ok... just to make sure \[\frac{dy}{dt} = \frac{e^y\sin^2(t)}{y\sec(t)}\]

OpenStudy (anonymous):

That is the correct equation

OpenStudy (anonymous):

\[ye^{-y}dy = \sin^{2}tcostdt\]

OpenStudy (cruffo):

looks like your on the right track with separation of variables. Move the e^y term over to get \[ye^{-y}dy = \frac{\sin^2(t)}{\sec(t)}\] simplify the right hand side to get what him1618 has...

OpenStudy (anonymous):

now for RHS take u = sin t

OpenStudy (cruffo):

integration by parts on the left hand side. U-substitution for the left hand side.

OpenStudy (anonymous):

What is the move that allows me to move e^y over

OpenStudy (anonymous):

divide by e^y both sides?

OpenStudy (cruffo):

multiply both sides by e^{-y}

OpenStudy (anonymous):

because 1/e^y = e^-y?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks folks

OpenStudy (anonymous):

no probs